3y2-3y-18
1/2m2n3-1/6mn2
Factorise plzz
1
Expert's answer
2012-10-16T09:33:51-0400
3y2-3y-18=3(y^2-y-6)=3(y+2)(y-3)

1/2m^2n^3-1/6mn^2=1/6 *m*n^2*(3mn-1)